# From Entanglement to Alignment: Representation Space Decomposition for Unsupervised Time Series Domain Adaptation

### v1

@@ Abstract
Domain shift poses a fundamental challenge in time series analysis, where models trained on source domain often fail dramatically when applied in target domain with different yet similar distributions. While current unsupervised domain adaptation (UDA) methods attempt to align cross-domain feature distributions, they typically treat features as indivisible entities, ignoring their intrinsic compositions that governs domain adaptation. We introduce DARSD, a novel UDA framework with theoretical explainability that explicitly realizes UDA tasks from the perspective of representation space decomposition. Our core insight is that effective domain adaptation requires not just alignment, but principled disentanglement of transferable knowledge from mixed representations. DARSD consists three synergistic components: (I) An adversarial learnable common invariant basis that projects original features into a domain-invariant subspace while preserving semantic content; (II) A prototypical pseudo-labeling mechanism that dynamically separates target features based on confidence, hindering error accumulation; (III) A hybrid contrastive optimization strategy that simultaneously enforces feature clustering and consistency while mitigating emerging distribution gaps. Comprehensive experiments conducted on four benchmark datasets (WISDM, HAR, HHAR, and MFD) demonstrate DARSD’s superiority against 12 UDA algorithms, achieving optimal performance in 35 out of 53 cross-domain scenarios.
@@ Introduction
In the context of rapid development of time series analysis, domain adaptation has emerged as a critical bottleneck of practical algorithm deployment across diverse environments (Zhang et al. 2024a; Fang et al. 2024). Consider a human activity recognition system trained on smartphone accelerometer data from young adults in a lab, its accuracy can drops sharply when used on smartwatches worn by elderly users in daily life (Zhao et al. 2025). This dramatic performance degradation stems from a fundamental issue: while the underlying patterns of human activities stay the same, like walking still exhibits periodic acceleration patterns, the data distributions shift significantly due to fluctuations in objects, environments, or perception (Zhao et al. 2021). Such domain shift scenarios are widespread across real-world applications ranging from industrial equipment monitoring (Cai et al. 2024a; An et al. 2025) to healthcare sensing (Ye and Wang 2024).
@@
Due to a lack of expert knowledge or high cost (Ozyurt, Feuerriegel, and Zhang 2023; Zhang et al. 2024b), the ground truth of target domain is not available during training in most scenarios. Unsupervised domain adaptation (UDA) has gained increasing attention, which trains a robust model using annotated source data with unlabeled target data, aiming to achieve good performance on the target domain (Zhang et al. 2024a; Westfechtel et al. 2024). Notably, source and target domain exhibit similar yet distinct distributions.
@@
Existing UDA approaches attempt to address domain shift through feature distribution alignment, they can be classified into three main categories. Adversarial training methods (Wang, Yang, and Yu 2024; Xie et al. 2025) learn domain-confusing representations by fooling discriminators (Lu et al. 2024). But this adversarial game often lacks supervision linked to downstream tasks, eliminating meaningful patterns alongside domain-specific artifacts. Metric learning approaches (Zhu et al. 2021; Zha et al. 2025) minimize statistical distances between source and target distributions (Lu et al. 2024), yet they impose an unrealistic assumption that entire features need be aligned, despite the fact that only part of components carry transferable knowledge. Self-supervised methods (Xu et al. 2025; He et al. 2023) depend on contrastive learning with pretext tasks, such as pseudo-labeling (Chen et al. 2021). However they run into a problem that label noise tends to build up progressively over the course of training. The core flaw underlying above approaches is that they treat features as monolithic and indivisible entities, implicitly hoping for invariance through distribution matching, overlooking the fact that effective domain adaptation demands a principled distinction between what ought to transfer and what ought not to. Though some methods have been proposed based on the insight of feature decomposition (Chen et al. 2024; Lu and Sun 2024), they rely on the implicit network structure or hidden variables.
@@
Our work is motivated by a key insight from representation learning: features possess internal structure. Domain shift primarily impacts certain dimensions of representations while leaving others unaltered. For instance, in accelerometer data, the temporal dynamics of walking remain largely consistent across devices (domain-invariant), whereas sensor noise characteristics and sampling artifacts vary significantly (domain-specific). This observation suggests a conceptual shift, instead of aligning entire feature distributions, we should explicitly decouple representations into orthogonal subspaces and selectively transfer only the invariant components. Such decomposition not only preserves semantic information but also provides interpretable insights into which patterns actually transfer across domains.
@@
Figure 1: The DARSD framework for time series domain adaptation. Features are first projected into a domain-invariant subspace (Adv-LCIB), then target features receive pseudo-labels with confidence evaluation (PPGCE), and finally all features undergo hybrid contrastive optimization for category-wise aggregation across domains. In hybrid contrastive optimization block, (1) Sup. Agg., (2) Self-sup. Consist., and (3) Anti-div. Reg. are the abbreviations of (1) Supervised Aggregation, (2) Self-supervised Consistence, and (3) Anti-divergence Regularization, respectively.
@@
[…]
@@
[…]
@@
Learnable Basis Construction: We introduce a learnable orthogonal matrix $B^{\textit{inv}}_{\textit{obs}}\in\mathbb{R}^{d\times m}$, spanning the domain-invariant subspace. For any given feature $f$, its domain-invariant representation is calculated as follow:
$w^{\textit{inv}}&=(B^{\textit{inv}}_{\textit{obs}})^{\top}f\\ &=(B^{\textit{inv}}_{\textit{obs}})^{\top}(f^{\textit{inv}}+f^{\textit{spe}})\\ &=(B^{\textit{inv}}_{\textit{obs}})^{\top}f^{\textit{inv}}=(B^{\textit{inv}}_{\textit{obs}})^{\top}B^{\textit{inv}}w^{\textit{inv}}\end{aligned}$ (2)
@@
To fully leverage target features, assigning reliable pseudo-labels to $\hat{F}_{t}$ is a viable solution, allowing them to join source features in supervised learning. This enables category-wise aggregation, where features from the same semantic class are clustered together regardless of domains, while those from different classes are separated. Even so, pseudo-labels vary in trustworthiness, that unreliable ones can introduce noise, degrading quality of clustering and disrupting decision boundaries. We need a mechanism to distinguish reliable from unreliable pseudo-labels and treat them differently during optimization.
@@
Anti-divergence regularization $\mathcal{L}_{\textit{anti}}$: A critical issue arises with the emerging distribution divergence caused by supervised aggregation and self-supervised consistency. Without proper coordination, these two optimization pathways may yield distinct feature distributions, violating the domain-invariant assumption and disrupting pseudo-label generation. Specifically, supervised aggregation produces cluster-wise distribution, while self-supervised consistency tends to generate uniform one (Cai et al. 2024b). To address this, we introduce an anti-divergence regularization that links distrusted target features to their semantically closest source counterparts. For each $\hat{f}_{i}^{\textit{dis}}$, we identify its nearest semantic neighbor in the source domain by maximizing cosine similarity:
@@
Table 1: Details of datasets.
Datasets #Subject #Channel #Length #Class Train Test
WISDM 30 3 128 6 56$\sim$246 30$\sim$132
HAR 30 9 128 6 211$\sim$286 85$\sim$123
HHAR 9 3 128 6 1066$\sim$1249 457$\sim$536
MFD 4 1 5120 3 1828 901
@@
(2) Metric learning approaches: These techniques focus on learning domain-invariant representations through various distance metrics and alignment strategies, encompassing DDC (Tzeng et al. 2014), DeepCORAL (Sun and Saenko 2016), CAN (Kang et al. 2019), HoMM (Chen et al. 2020), MMDA (Rahman et al. 2020), AdvSKM (Liu and Xue 2021), and DSAN (Zhu et al. 2021), and DARSD (ours);
@@
Table 2: Macro-F1 score of UDA methods on WISDM, HAR, HHAR, and MFD datasets. Each result is the mean over ten random initializations. Src and Trg denote the source and target subsets. The best two results are highlighted by 1st and 2nd.
Src $\to$ Trg DDC DeepCORAL VRADA CDAN CAN CoDATS HoMM MMDA AdvSKM DSAN CLUDA CADT DARSD (ours)
WISDM 0 $\to$ 13 0.319 0.608 0.441 0.381 0.319 0.347 0.577 0.288 0.563 0.577 0.409 0.614 0.625
WISDM 1 $\to$ 24 0.875 0.746 0.750 0.360 0.515 0.500 0.418 0.360 0.304 0.676 0.812 0.381 1.000
WISDM 2 $\to$ 28 0.669 0.726 0.688 0.644 0.610 0.688 0.691 0.677 0.742 0.654 0.788 0.693 0.822
WISDM 4 $\to$ 11 0.309 0.585 0.264 0.231 0.146 0.226 0.207 0.226 0.211 0.578 0.614 0.448 0.712
WISDM 5 $\to$ 12 0.484 0.511 0.573 0.094 0.521 0.473 0.508 0.086 0.515 0.405 0.518 0.653 0.577
WISDM 7 $\to$ 2 0.496 0.490 0.399 0.543 0.490 0.494 0.494 0.459 0.476 0.481 0.576 0.632 0.667
WISDM 7 $\to$ 26 0.412 0.396 0.308 0.344 0.395 0.405 0.406 0.385 0.416 0.401 0.403 0.447 0.456
WISDM 10 $\to$ 14 0.541 0.337 0.591 0.266 0.608 0.477 0.667 0.469 0.471 0.337 0.510 0.599 0.691
WISDM 10 $\to$ 21 0.694 0.250 0.125 0.118 0.250 0.390 0.313 0.390 0.694 0.106 0.250 0.518 0.392
WISDM 12 $\to$ 7 0.632 0.486 0.437 0.546 0.636 0.612 0.442 0.539 0.655 0.574 0.678 0.620 0.834
WISDM 12 $\to$ 19 0.396 0.317 0.410 0.298 0.508 0.456 0.281 0.233 0.510 0.518 0.532 0.518 0.883
WISDM 15 $\to$ 29 0.299 0.265 0.391 0.238 0.372 0.136 0.208 0.133 0.299 0.182 0.482 0.441 0.529
WISDM 18 $\to$ 20 0.383 0.379 0.578 0.600 0.389 0.427 0.421 0.280 0.348 0.268 0.673 0.702 0.658
WISDM 19 $\to$ 2 0.459 0.501 0.615 0.312 0.327 0.403 0.522 0.306 0.460 0.428 0.458 0.555 0.664
WISDM 21 $\to$ 16 0.537 0.669 0.542 0.404 0.641 0.410 0.669 0.310 0.523 0.568 0.595 0.850 0.787
WISDM 25 $\to$ 22 0.273 0.241 0.506 0.401 0.241 0.200 0.307 0.159 0.309 0.341 0.375 0.500 0.610
WISDM 26 $\to$ 2 0.414 0.618 0.517 0.404 0.362 0.598 0.519 0.453 0.463 0.424 0.701 0.610 0.731
WISDM 28 $\to$ 2 0.484 0.495 0.473 0.400 0.412 0.492 0.511 0.430 0.484 0.451 0.710 0.606 0.656
WISDM 28 $\to$ 20 0.571 0.620 0.672 0.605 0.655 0.578 0.699 0.537 0.557 0.615 0.703 0.726 0.728
@@
HAR 2 $\to$ 11 0.372 0.670 0.680 0.358 0.644 0.434 0.545 0.374 0.385 0.428 0.644 0.782 0.717
HAR 3 $\to$ 20 0.671 0.853 0.847 0.549 0.769 0.852 0.828 0.757 0.860 0.784 0.968 0.864 0.907
HAR 5 $\to$ 18 0.552 0.595 0.715 0.491 0.636 0.626 0.727 0.462 0.668 0.663 0.659 0.710 0.745
HAR 7 $\to$ 13 0.511 0.436 0.680 0.442 0.530 0.505 0.611 0.388 0.533 0.639 0.884 0.886 0.681
HAR 9 $\to$ 18 0.453 0.665 0.715 0.357 0.362 0.389 0.508 0.319 0.461 0.590 0.610 0.653 0.846
HAR 12 $\to$ 16 0.446 0.663 0.689 0.655 0.500 0.432 0.678 0.385 0.446 0.561 0.582 0.585 0.832
HAR 13 $\to$ 19 0.774 0.716 0.797 0.548 0.619 0.852 0.852 0.582 0.828 0.906 0.906 0.811 0.864
HAR 18 $\to$ 21 0.439 0.760 0.697 0.450 0.449 0.857 0.803 0.380 0.439 0.588 0.920 0.929 0.948
HAR 20 $\to$ 6 0.643 0.776 0.619 0.407 0.593 0.659 0.935 0.407 0.586 0.588 0.959 0.782 0.967
HAR 23 $\to$ 13 0.398 0.586 0.683 0.298 0.509 0.439 0.539 0.518 0.401 0.723 0.720 0.743 0.764
HAR 24 $\to$ 12 0.724 0.765 0.658 0.758 0.872 0.709 0.893 0.448 0.736 0.826 0.866 0.768 0.887
Avg Rank 10.2 6.55 5.55 11.0 8.45 8.18 4.82 12.0 8.36 6.36 3.73 3.64 1.64
HHAR 0 $\to$ 2 0.605 0.569 0.536 0.611 0.598 0.598 0.627 0.612 0.628 0.415 0.710 0.585 0.651
HHAR 0 $\to$ 4 0.266 0.261 0.284 0.264 0.257 0.253 0.317 0.214 0.341 0.443 0.589 0.410 0.619
HHAR 1 $\to$ 6 0.678 0.725 0.702 0.727 0.621 0.696 0.726 0.693 0.662 0.696 0.858 0.782 0.861
HHAR 2 $\to$ 3 0.297 0.437 0.380 0.338 0.408 0.444 0.303 0.253 0.393 0.598 0.582 0.646 0.692
HHAR 2 $\to$ 4 0.231 0.305 0.415 0.431 0.294 0.320 0.230 0.192 0.219 0.143 0.526 0.464 0.480
HHAR 4 $\to$ 0 0.175 0.249 0.243 0.273 0.165 0.222 0.179 0.162 0.163 0.116 0.352 0.304 0.483
HHAR 4 $\to$ 1 0.456 0.461 0.545 0.667 0.523 0.469 0.607 0.517 0.466 0.488 0.751 0.646 0.796
HHAR 5 $\to$ 2 0.311 0.330 0.327 0.324 0.338 0.209 0.335 0.278 0.229 0.232 0.299 0.318 0.340
HHAR 7 $\to$ 0 0.194 0.198 0.169 0.454 0.178 0.282 0.232 0.095 0.189 0.247 0.438 0.304 0.527
HHAR 7 $\to$ 2 0.233 0.295 0.525 0.333 0.250 0.217 0.258 0.235 0.203 0.179 0.332 0.390 0.392
HHAR 7 $\to$ 4 0.577 0.608 0.658 0.344 0.736 0.489 0.750 0.539 0.543 0.522 0.604 0.850 0.813
HHAR 7 $\to$ 5 0.488 0.496 0.529 0.480 0.546 0.374 0.323 0.283 0.487 0.649 0.626 0.645 0.680
HHAR 8 $\to$ 6 0.568 0.728 0.723 0.334 0.765 0.701 0.745 0.514 0.734 0.467 0.837 0.773 0.844
Avg Rank 9.38 7.31 6.85 6.77 7.54 9.00 6.62 10.7 9.15 9.00 3.38 3.85 1.31
MFD 1 $\to$ 0 0.500 0.514 0.427 0.555 0.507 0.614 0.506 0.515 0.500 0.484 0.494 0.709 0.663
MFD 1 $\to$ 2 0.749 0.704 0.500 0.881 0.572 0.919 0.747 0.773 0.734 0.732 0.574 0.680 0.776
MFD 0 $\to$ 1 0.524 0.504 0.455 0.599 0.518 0.607 0.509 0.593 0.468 0.374 0.510 0.515 0.722
MFD 0 $\to$ 3 0.531 0.513 0.463 0.598 0.468 0.614 0.504 0.644 0.493 0.439 0.605 0.546 0.716
MFD 1 $\to$ 3 0.940 0.892 0.521 0.987 0.625 0.994 0.940 0.997 0.930 0.996 0.900 0.765 0.997
MFD 2 $\to$ 1 0.893 0.714 0.493 0.997 0.587 0.994 0.849 0.971 0.887 0.993 0.559 0.613 0.953
MFD 0 $\to$ 2 0.762 0.490 0.494 0.772 0.477 0.502 0.540 0.701 0.508 0.426 0.689 0.555 0.770
MFD 3 $\to$ 2 0.737 0.695 0.485 0.899 0.564 0.906 0.729 0.783 0.728 0.846 0.790 0.686 0.813
MFD 2 $\to$ 0 0.659 0.609 0.445 0.689 0.489 0.669 0.613 0.594 0.562 0.627 0.628 0.637 0.792
MFD 3 $\to$ 0 0.573 0.527 0.474 0.552 0.471 0.679 0.521 0.526 0.507 0.511 0.626 0.736 0.756
Avg Rank 5.70 8.90 12.4 3.10 10.8 3.10 7.80 4.80 9.10 8.50 7.50 6.90 2.10
Table 3: Ablation analysis of DARSD. Specifically, the contributions of LCIB, $\mathcal{L}_{\textit{adv}}$, $\mathcal{L}_{\textit{self}}$, and $\mathcal{L}_{\textit{anti}}$ are verified. The check mark ($\checkmark$) denotes the activated components. We bold the best results and underlined the runner-ups.
@@
LCIB $\mathcal{L}_{\textit{adv}}$ $\mathcal{L}_{\textit{sup}}$ $\mathcal{L}_{\textit{self}}$ $\mathcal{L}_{\textit{anti}}$ WISDM 121912$\to$ 1912 → 19 HAR 20620$\to$ 620 → 6 HHAR 868$\to$ 68 → 6 WISDM 2282$\to$ 282 → 28 HHAR 747$\to$ 47 → 4
Macro-F1 Acc Macro-F1 Acc Macro-F1 Acc Macro-F1 Acc Macro-F1 Acc
@@
The comparative analysis illuminates advantages of our DARSD over existing paradigms. Adversarial methods ($e.g.$, CDAN and VRADA) exhibit inconsistent performance, as they risk eliminating not only domain-specific artifacts while also also valid semantic information. Metric learning approaches ($e.g.$, HoMM and CADT) demonstrate moderate consistency. While the self-supervision method CLUDA performs strongly, relying on gradually and slowly contrastive learning and adversarial training. In contrast, our DARSD framework provides more direct control over trade-off between invariance and discriminability, with faster convergence. The Convergence Efficiency of DARSD and CLUDA are compared in Appendix D.3. To better characterize performance of DARSD, we conduct Representation Visualization of source and target features in Appendix D.2.
@@
Synergistic Effects: The great performance gap between any individual element and the full model demonstrates that DARSD’s effectiveness arises from the synergistic interaction of all components, rather than merely the sum of their individual contributions. Each element addresses specific challenges, $i.e.$, LCIB handles representation decomposition, $\mathcal{L}_{\textit{adv}}$ ensures domain invariance, $\mathcal{L}_{\textit{sup}}$ aggregates same-category features, $\mathcal{L}_{\textit{self}}$ enhances distrusted target feature quality, and $\mathcal{L}_{\textit{anti}}$ maintains distribution coherence.
Notations: Due to space constraints, Proof of Representation Space Decomposition, Representation Visualization, Sensitivity Analysis of Hyperparameters, Convergence Efficiency, and Related Works are provided in the Appendix.

### v2

@@ Abstract
Domain shift poses a fundamental challenge in time series analysis, where models trained on source domain often fail dramatically when applied in target domain with different yet similar distributions. While current unsupervised domain adaptation (UDA) methods attempt to align cross-domain feature distributions, they typically treat features as indivisible entities, ignoring their intrinsic compositions that govern domain adaptation. We introduce DARSD, a novel UDA framework with theoretical explainability that explicitly realizes UDA tasks from the perspective of representation space decomposition. Our core insight is that effective domain adaptation requires not just alignment, but principled disentanglement of transferable knowledge from mixed representations. DARSD consists of three synergistic components: (I) An adversarial learnable common invariant basis that projects original features into a domain-invariant subspace while preserving semantic content; (II) A prototypical pseudo-labeling mechanism that dynamically separates target features based on confidence, hindering error accumulation; (III) A hybrid contrastive optimization strategy that simultaneously enforces feature clustering and consistency while mitigating emerging distribution gaps. Comprehensive experiments conducted on four benchmarks (WISDM, HAR, HHAR, and MFD) demonstrate DARSD’s superiority against 12 UDA algorithms, achieving optimal performance in 35 out of 53 scenarios and ranking first across all benchmarks.
@@ Introduction
In the context of rapid development of time series analysis, domain adaptation has emerged as a critical bottleneck of practical algorithm deployment across diverse environments (Zhang et al. 2024a; Fang et al. 2024). Consider a human activity recognition system trained on smartphone accelerometer data from young adults in a lab, its accuracy can drop sharply when used on smartwatches worn by elderly users in daily life (Zhao et al. 2025). This dramatic performance degradation stems from a fundamental issue: while the underlying patterns of human activities stay the same, like walking still exhibits periodic acceleration patterns, the data distributions shift significantly due to fluctuations in objects, environments, or perception (Zhao et al. 2021). Such domain shift scenarios are widespread across real-world applications ranging from industrial equipment monitoring (Cai et al. 2024a; An et al. 2025) to healthcare sensing (Ye and Wang 2024).
@@
Due to a lack of expert knowledge or high cost (Ozyurt, Feuerriegel, and Zhang 2023; Zhang et al. 2024b), the ground truth of target domain is not available during training in most scenarios. Unsupervised domain adaptation (UDA) has gained increasing attention, which trains a robust model using annotated source data with unlabeled target data, aiming to achieve good performance on the latter (Zhang et al. 2024a; Westfechtel et al. 2024). Notably, source and target domain exhibit similar yet distinct distributions.
Figure 1: Visual description of representation space decomposition. We disentangle the basis of representation space into invariant one $B^{\textit{inv}}$ and specific ones related to domains $B^{\textit{spe}}_{s}$ and $B^{\textit{spe}}_{t}$, aiming to achieve discriminative clustering in domain-invariant subspace spanned by $B^{\textit{inv}}$.
@@
Existing UDA approaches attempt to address domain shift through feature distribution alignment, they can be classified into three main categories. Adversarial training methods (Wang, Yang, and Yu 2024; Xie et al. 2025) learn domain-confusing representations by fooling discriminators (Lu et al. 2024). But this adversarial game often lacks supervision linked to downstream tasks, eliminating meaningful patterns alongside domain-specific artifacts. Metric learning approaches (Zhu et al. 2021; Zha et al. 2025) minimize statistical distances between source and target distributions (Lu et al. 2024), yet they impose an unrealistic assumption that entire features need be aligned, despite the fact that only part of components carry transferable knowledge. Self-supervised methods (Xu et al. 2025; He et al. 2023) depend on contrastive learning with pretext tasks, such as pseudo-labeling (Chen et al. 2021). However, they run into a problem that label noise tends to build up progressively over the course of training. The core flaw underlying above approaches is that they treat features as monolithic and indivisible entities, implicitly hoping for invariance through distribution matching, overlooking the fact that effective domain adaptation demands a principled distinction between what ought to transfer and what ought not to. Though several methods have been proposed based on the insight of feature decomposition, they rely on the implicit network structure (Chen et al. 2024; Qu et al. 2024) or causal hidden variables (Lu and Sun 2024).
@@
Our work is motivated by a key insight from representation learning: features possess internal structure, as depicted in Fig. 1. Domain shift primarily impacts certain dimensions of representations while leaving others unaltered. For instance, in accelerometer data, the temporal dynamics of walking remain largely consistent across devices (domain-invariant), whereas sensor noise characteristics and sampling artifacts vary significantly (domain-specific). This observation suggests a conceptual shift, instead of aligning entire feature distributions, we should explicitly decouple representations into orthogonal subspaces and selectively transfer only the invariant components. Such decomposition not only preserves semantic information but also provides interpretable insights into which patterns actually transfer across domains.
@@
Figure 2: The DARSD framework for time series domain adaptation. Features are first projected into a domain-invariant subspace (Adv-LCIB), then target features receive pseudo-labels with confidence evaluation (PPGCE), and finally all features undergo hybrid contrastive optimization for category-wise aggregation across domains. In hybrid contrastive optimization block, (1) Sup. Agg., (2) Self-sup. Consist., (3) Anti-div. Reg., and (4) ProxSrch are the abbreviations of (1) Supervised Aggregation, (2) Self-supervised Consistence, (3) Anti-divergence Regularization, and (4) Proximity Searcher, respectively.
@@
Our objective is to train a robust feature extractor $\textit{FE}(\cdot)$ by leveraging annotated $\mathcal{S}$ and unlabeled $\mathcal{T}$. This extractor will derive a domain-invariant feature set $F=\{F_{s}\cup F_{t}\}^{n_{s}+n_{t}}$ from $\mathcal{S}$ and $\mathcal{T}$. Subsequently, we fine-tune a classifier $\textit{CLF}(\cdot)$ using $\{(f^{s}_{i},y^{s}_{i})\}_{i=1}^{n_{s}}$ to predict the label set $Y_{t}=\{y^{t}_{i}\}_{i=1}^{n_{t}}$ for $\mathcal{T}$ based on the feature set $F_{t}=\{f^{t}_{i}\}_{i=1}^{n_{t}}$.
@@
The detailed pre-training procedure of our DARSD framework is illustrated in Fig. 2. DARSD comprises a shared feature extractor $\textit{FE}(\cdot)$, an Adv-LCIB module, and a hybrid contrastive optimization mechanism integrated with the PPGCE module.
@@
Learnable Basis Construction: We introduce a learnable orthogonal matrix $B^{\textit{inv}}_{\textit{obs}}\in\mathbb{R}^{d\times m}$, spanning the domain-invariant subspace. For any given feature $f$, its domain-invariant representation is calculated as follows:
$\begin{aligned} w^{\textit{inv}}&=(B^{\textit{inv}}_{\textit{obs}})^{\top}f\\ &=(B^{\textit{inv}}_{\textit{obs}})^{\top}(f^{\textit{inv}}+f^{\textit{spe}})\\ &=(B^{\textit{inv}}_{\textit{obs}})^{\top}f^{\textit{inv}}=(B^{\textit{inv}}_{\textit{obs}})^{\top}B^{\textit{inv}}w^{\textit{inv}}\end{aligned}$ (2)
@@
To fully leverage target features, assigning reliable pseudo-labels to $\hat{F}_{t}$ is a viable solution, allowing them to join source features in supervised learning. This enables category-wise aggregation, where features from the same semantic class are clustered together regardless of domains, while those from different classes are separated. Even so, pseudo-labels vary in trustworthiness, that unreliable ones can introduce noise, degrade the quality of clustering, and disrupt decision boundaries. We need a mechanism to distinguish reliable from unreliable pseudo-labels and treat them differently during optimization.
@@
Anti-divergence regularization $\mathcal{L}_{\textit{anti}}$: A critical issue arises with the emerging distribution divergence caused by supervised aggregation and self-supervised consistency. Without proper coordination, these two optimization pathways may yield distinct feature distributions, violating the domain-invariant assumption and disrupting pseudo-label generation. Specifically, supervised aggregation produces cluster-wise distribution, while self-supervised consistency tends to generate uniform one (Cai et al. 2024b). To address this, we introduce an anti-divergence regularization that links distrusted target features to their semantically closest source counterparts. For each $\hat{f}_{i}^{\textit{dis}}$, we identify its nearest semantic neighbor in the source domain by proximity searcher operator $\textit{PS}(\cdot)$:
@@
Datasets $D$ $T$ $n_{c}$ Subject Train Test
WISDM 3 128 6 30 1350 720
HAR 9 128 6 30 2300 990
HHAR 3 128 6 9 12716 5218
MFD 1 5120 3 4 7312 3604
Table 1: Characteristics of benchmarks.
@@
(2) Metric learning approaches: These techniques focus on learning domain-invariant representations through various distance metrics and alignment strategies, encompassing DDC (Tzeng et al. 2014), D-CORAL (Sun and Saenko 2016), CAN (Kang et al. 2019), HoMM (Chen et al. 2020), MMDA (Rahman et al. 2020), AdvSKM (Liu and Xue 2021), and DSAN (Zhu et al. 2021), and DARSD (ours);
@@
Src $\to$ Trg DDC D-CORAL VRADA CDAN CAN CoDATS HoMM MMDA AdvSKM DSAN CLUDA CADT DARSD
WISDM; Out of 19 scenarios, DARSD achieves the 15 optimal results and 3 runner-up positions.
0 $\to$ 13 0.319 0.608 0.441 0.381 0.319 0.347 0.577 0.288 0.563 0.577 0.409 0.614 0.625
1 $\to$ 24 0.875 0.746 0.750 0.360 0.515 0.500 0.418 0.360 0.304 0.676 0.812 0.381 1.000
2 $\to$ 28 0.669 0.726 0.688 0.644 0.610 0.688 0.691 0.677 0.742 0.654 0.788 0.693 0.822
4 $\to$ 11 0.309 0.585 0.264 0.231 0.146 0.226 0.207 0.226 0.211 0.578 0.614 0.448 0.712
5 $\to$ 12 0.484 0.511 0.573 0.094 0.521 0.473 0.508 0.086 0.515 0.405 0.518 0.653 0.577
7 $\to$ 2 0.496 0.490 0.399 0.543 0.490 0.494 0.494 0.459 0.476 0.481 0.576 0.632 0.667
7 $\to$ 26 0.412 0.396 0.308 0.344 0.395 0.405 0.406 0.385 0.416 0.401 0.403 0.447 0.456
@@
HAR; Out of 11 scenarios, DARSD achieves the 6 optimal results and 4 runner-up positions.
2 $\to$ 11 0.372 0.670 0.680 0.358 0.644 0.434 0.545 0.374 0.385 0.428 0.644 0.782 0.717
3 $\to$ 20 0.671 0.853 0.847 0.549 0.769 0.852 0.828 0.757 0.860 0.784 0.968 0.864 0.907
5 $\to$ 18 0.552 0.595 0.715 0.491 0.636 0.626 0.727 0.462 0.668 0.663 0.659 0.710 0.745
7 $\to$ 13 0.511 0.436 0.680 0.442 0.530 0.505 0.611 0.388 0.533 0.639 0.884 0.886 0.681
9 $\to$ 18 0.453 0.665 0.715 0.357 0.362 0.389 0.508 0.319 0.461 0.590 0.610 0.653 0.846
12 $\to$ 16 0.446 0.663 0.689 0.655 0.500 0.432 0.678 0.385 0.446 0.561 0.582 0.585 0.832
13 $\to$ 19 0.774 0.716 0.797 0.548 0.619 0.852 0.852 0.582 0.828 0.906 0.906 0.811 0.864
Avg. Rank 10.2 6.55 5.55 11.0 8.45 8.18 4.82 12.0 8.36 6.36 3.73 3.64 1.64
HHAR; Out of 13 scenarios, DARSD achieves the 9 optimal results and 4 runner-up positions.
0 $\to$ 2 0.605 0.569 0.536 0.611 0.598 0.598 0.627 0.612 0.628 0.415 0.710 0.585 0.651
0 $\to$ 4 0.266 0.261 0.284 0.264 0.257 0.253 0.317 0.214 0.341 0.443 0.589 0.410 0.619
1 $\to$ 6 0.678 0.725 0.702 0.727 0.621 0.696 0.726 0.693 0.662 0.696 0.858 0.782 0.861
2 $\to$ 3 0.297 0.437 0.380 0.338 0.408 0.444 0.303 0.253 0.393 0.598 0.582 0.646 0.692
2 $\to$ 4 0.231 0.305 0.415 0.431 0.294 0.320 0.230 0.192 0.219 0.143 0.526 0.464 0.480
4 $\to$ 0 0.175 0.249 0.243 0.273 0.165 0.222 0.179 0.162 0.163 0.116 0.352 0.304 0.483
4 $\to$ 1 0.456 0.461 0.545 0.667 0.523 0.469 0.607 0.517 0.466 0.488 0.751 0.646 0.796
Avg. Rank 9.38 7.31 6.85 6.77 7.54 9.00 6.62 10.7 9.15 9.00 3.38 3.85 1.31
MFD; Out of 10 scenarios, DARSD achieves the 5 optimal results and 2 runner-up positions.
0 $\to$ 3 0.531 0.513 0.463 0.598 0.468 0.614 0.504 0.644 0.493 0.439 0.605 0.546 0.716
1 $\to$ 3 0.940 0.892 0.521 0.987 0.625 0.994 0.940 0.997 0.930 0.996 0.900 0.765 0.997
2 $\to$ 1 0.893 0.714 0.493 0.997 0.587 0.994 0.849 0.971 0.887 0.993 0.559 0.613 0.953
0 $\to$ 2 0.762 0.490 0.494 0.772 0.477 0.502 0.540 0.701 0.508 0.426 0.689 0.555 0.770
3 $\to$ 2 0.737 0.695 0.485 0.899 0.564 0.906 0.729 0.783 0.728 0.846 0.790 0.686 0.813
2 $\to$ 0 0.659 0.609 0.445 0.689 0.489 0.669 0.613 0.594 0.562 0.627 0.628 0.637 0.792
3 $\to$ 0 0.573 0.527 0.474 0.552 0.471 0.679 0.521 0.526 0.507 0.511 0.626 0.736 0.756
Avg. Rank 5.70 8.90 12.4 3.10 10.8 3.10 7.80 4.80 9.10 8.50 7.50 6.90 2.10
Table 2: Partial Macro-F1 results on WISDM, HAR, HHAR, and MFD benchmarks. Each result is the mean over ten random initializations. Src and Trg denote the source and target subsets. The best two results are highlighted by 1st and 2nd. The remaining results appear in Appendix Table 2.
@@
LCIB $\mathcal{L}_{\textit{adv}}$ $\mathcal{L}_{\textit{sup}}$ $\mathcal{L}_{\textit{self}}$ $\mathcal{L}_{\textit{anti}}$ WISDM $12\to 19$ HAR $20\to 6$ HHAR $8\to 6$ WISDM $2\to 28$ HHAR $7\to 4$
MF1 Acc MF1 Acc MF1 Acc MF1 Acc MF1 Acc
@@
Table 3: Ablation analysis of DARSD. Specifically, the contributions of LCIB, $\mathcal{L}_{\textit{adv}}$, $\mathcal{L}_{\textit{self}}$, and $\mathcal{L}_{\textit{anti}}$ are verified. The check mark ($\checkmark$) denotes the activated components. We bold the best results and underlined the runner-up positions.
@@
The comparative analysis illuminates advantages of our DARSD over existing paradigms. Adversarial methods ($e.g.$, CDAN and VRADA) exhibit inconsistent performance, as they risk eliminating not only domain-specific artifacts while also also valid semantic information. Metric learning approaches ($e.g.$, HoMM and CADT) demonstrate moderate consistency. While the self-supervision method CLUDA performs strongly, relying on gradually and slowly contrastive learning and adversarial training. In contrast, our DARSD framework provides more direct control over trade-off between invariance and discriminability, with faster convergence. The Convergence Efficiency of DARSD and CLUDA is compared in Appendix D.4. To better characterize performance of DARSD, we conduct Representation Visualization of source and target features in Appendix D.3.
@@
Synergistic Effects: The significant performance gap between any individual component and the full DARSD model demonstrates that its effectiveness arises from component synergy, not mere aggregation.
Notations: Due to space constraints, A. Notation, B. Proof of Representation Space Decomposition, C. Detailed Benchmark Description, D.1. Macro-F1 Results, D.2. Hyperparameter Sensitivity Analysis, D.3. Representation Visualization, D.4. Convergence Efficiency, and E. Related Works are provided in Appendix.